# Annamaria Porreca Fabrizio Maturo Viviana Ventre

### v1

@@ Abstract
Social network analysis (SNA) helps us understand the relationships and interactions between individuals, groups, organisations, or other social entities. In SNA, ties are generally binary or weighted based on their strength. Nonetheless, when actors are individuals, the relationships between actors are often imprecise and identifying them with simple scalars leads to information loss. Social relationships are often vague in real life. Despite many classical social network techniques contemplate the use of weighted links, these approaches do not align with the original philosophy of fuzzy logic, which instead aims to preserve the vagueness inherent in human language and real life. Dealing with imprecise ties and introducing fuzziness in the definition of relationships requires an extension of social network analysis to fuzzy numbers instead of crisp values. The mathematical formalisation for this generalisation needs to extend classical centrality indices and operations to fuzzy numbers. For this reason, this paper proposes a generalisation of the so-called Fuzzy Social Network Analysis (FSNA) to the context of imprecise relationships among actors. The article shows the theory and application of real data collected through a fascinating mouse tracking technique to study the fuzzy relationships in a collaboration network among the members of a University department.
keywords:
Fuzzy Social Network , Fuzzy Centrality Indices , Vague Relationships , Uncertainty , Fuzzy Numbers , University Department Collaboration Network.
MSC:
62A86 , 94D05 , 91D30
[…]
[inst1]organization=Department of Medical, Oral and Biotechnologies Science, University G. d’Annunzio of Chieti-Pescara, addressline=Via dei Vestini, 31, city=Chieti, postcode=66100, country=Italy
[…]
[inst2]organization=Faculty of Technological and Innovation Sciences, Universitas Mercatorum, addressline=Piazza Mattei, 10, city=Rome, postcode=00186, country=Italy
[…]
[inst3]organization=Department of Mathematics and Physics, University of Campania Luigi Vanvitelli, addressline=Viale Abramo Lincoln, 5, city=Caserta, postcode=81100, country=Italy
@@ 1 Introduction
Social network analysis (SNA) helps us understand the relationships and interactions between individuals, groups, organisations, or other social entities [18, 31, 29, 40, 26]. Each type of social structure can be represented with a graph, in which the social entities are represented as nodes or vertices in a network, and their relationships are represented as edges or ties. SNA examines relationships between entities and how information, resources, or influence flows through them. It identifies patterns, structures, and dynamics of social systems. Ties in a network are generally binary based on the presence or absence of the link or weighted based on their strength. Each connection can be identified by a number in $[0,1]$ or $\{0,1\}$, respectively. Nonetheless, when the actors of a social network are individuals, the relationships are very complex and measuring them using a single number leads to a loss of information. In other words, relationships are inherently imprecise in real life, and thus, social network analysis should consider vagueness. For example, let’s consider a social network like Facebook. The nodes may represent people registered on the network, and the links can be the connections between pairs of individuals. Typically, in a social network of this type, these connections are based on the presence or absence of ”friendship”. We must consider whether being friends with someone on Facebook should have the same significance for all connections. It’s evident that not all friendships are of equal weight and intensity; the level of interaction determines the closeness of a relationship rather than merely being connected online. However, this issue applies to social networks and real-life relationships because connections are fuzzy and real-life phenomena are not black or white but often grey [45, 47, 46]. What does it mean to be a friend? Is it being a close friend, an acquaintance, or just a casual acquaintance? What does being close collaborators or barely knowing each other in a university department mean?
One way to address this issue is using a fuzzy network analysis approach. Despite this, fuzzy network analysis has mainly been studied by introducing weighted ties within the network [see, e.g., 20, 8]. This approach of weighing the links between nodes differs from the original philosophy of fuzzy logic, which instead aims to preserve the vagueness inherent in human language. Assigning a link a value in $[0,1]$ or $\{0,1\}$ can sometimes be a simplification that sacrifices information for the sake of convenience. Translating vague concepts (such as close collaborator, occasional collaborator, real friend, and similar) necessarily results in a loss of information regarding the inherently imprecise nature of these expressions. Hence, considering a link as ”fuzzy” merely because it is represented by a number in the interval $[0,1]$ does not align with the basic principles of fuzzy logic, which begins and thrives where opposites coexist. In other words, fuzzy logic deals with situations where A and non-A coexist simultaneously with a certain degree of truth [45, 47, 46, 41]. Dealing with imprecise ties and introducing fuzziness in the definition of relationships needs an extension of classical social network analysis to fuzzy numbers instead of crisp values. The mathematical formalisation of this generalisation requires rewriting the classical centrality indices into fuzzy centrality indices involving fuzzy relationships. In other words, this research’s primary motivation and basic idea is that relationships between nodes should be characterised by functions rather than scalars to preserve the information about the vagueness of ties. The great advantage of this approach is the ability to create social networks and centrality indices that preserve information about the vagueness and imprecision of ties because they are inherent in human nature, human language, and interpersonal relationships. In statistics, it is known that valuable information is contained within variability. Similarly, in social networks, true information lies in understanding the vagueness of relationships.
@@
The article shows the theory of Fuzzy Centrality Measures (FCM) in Social Network Analysis and their application to real data. The dataset adopted in this study is collected through a fascinating technique that includes mouse tracking to capture the vagueness of the relationships among the members of a University department. The first part of the work introduces the preliminaries necessary for formally understanding the proposed indices. Subsequently, based on the idea mentioned above, new centrality indices are proposed: fuzzy degree centrality, fuzzy out-degree centrality index, total fuzzy degree centrality index, fuzzy betweenness centrality, fuzzy closeness centrality, and fuzzy out-closeness centrality. The application Section proposes a fascinating investigation based on a University department collaboration network. The paper ends with a discussion and conclusions.
2 A brief introduction to Graph Theory (GT)
A graph $G=(V,E)$ is an abstract object formed by a set $V$ of vertices (nodes) and a set $E$ of edges (links) that join pairs of vertices. The vertex set and edge set of a graph $G$ are denoted by $V(G)$ and $E(G)$, respectively. The cardinality of $V$ is usually denoted by $n$, and the cardinality of $E$ by $m$. If an edge joins two vertices, they are adjacent, and we call them neighbors. Graphs can be undirected or directed. In undirected graphs, the order of the end vertices of an edge does not indicate a specific direction between them; the relationship is symmetric. In directed graphs, each directed edge (arc) has an origin (tail) and a destination (head). An edge with origin $u\in V$ and destination $v\in V$ is represented by an ordered pair $(u,v)$. In shorthand notation, an edge $\{u,v\}$ or $(u,v)$ can also be denoted by $uv$. In a directed graph, $uv$ is short for $(u,v)$, while in an undirected graph, $uv$ and $vu$ are the same and both stand for $\{u,v\}$. Edge weights are numerical values associated with edges or vertices in a graph $G=(V,E)$. Here, we focus on edge weights. These weights are often represented by a function $w:E\to\mathbb{R}$, assigning each edge $e\in E$ a weight $w(e)$. Depending on the application, edge weights can represent properties such as cost, capacity, strength of interaction, or similarity.
A walk from $v_{0}$ to $v_{k}$ in a graph $G=(V,E)$ is an alternating sequence $v_{0},e_{1},v_{1},e_{2},v_{2},$
$\ldots,v_{k-1},e_{k},v_{k}$ of vertices and edges, where $e_{i}=\{v_{i-1},v_{i}\}$ in the undirected case and $e_{i}=(v_{i-1},v_{i})$ in the directed case. The length of the walk is defined as the number of edges on the walk. The walk is called a path if $e_{i}\neq e_{j}$ for $i\neq j$, and a path is a simple path if $v_{i}\neq v_{j}$ for $i\neq j$. A path with $v_{0}=v_{k}$ is a cycle. A cycle is a simple cycle if $v_{i}\neq v_{j}$ for $0\leq i<j\leq(k-1)$. For a path $p$ in a graph $G=(V,E)$ with edge weights $w$, the weight of the path, denoted by $w(p)$, is defined as the sum of the weights of the edges on $p$. A path from $u$ to $v$ in $G$ is the shortest path (with respect to $w$) if its weight is the smallest possible among all paths from $u$ to $v$. The length of the shortest path from $u$ to $v$, also called the shortest-path distance between $u$ and $v$, is denoted by $d(u,v)$.
The single-source shortest paths problem is defined as follows: Given a graph $G=(V,E)$ with edge weights $w:E\to\mathbb{R}$ and a vertex $v\in V$ (the source), compute the shortest paths from $v$ to all other vertices in the graph. The problem is only well-defined if the graph does not contain a negative weight cycle. If the edge weights are non-negative, the shortest paths problem can be solved in time $O(m+n\log n)$ using an efficient implementation of Dijkstra’s algorithm [15]. If the edge weights are arbitrary, the Bellman-Ford algorithm [4] uses time $O(mn)$ to detect a cycle of negative length or, if no such cycle exists, solve the problem. For the case of unit edge weights, the Bellman-Ford algorithm solves the problem in linear time $O(n+m)$. In the all-pairs shortest paths problem, one is given a graph $G=(V,E)$ with edge weights $w:E\to\mathbb{R}$ and wants to compute the shortest-path distances for all pairs of nodes. Provided that $G$ does not contain a cycle of negative length, this problem can be solved by the Floyd-Warshall algorithm [17] in time $O(n^{3})$, or by running the shortest paths’ problem computations in time $O(nm+n^{2}\log n)$.
Adjacency and incidence matrices are tools used in graph theory to represent and analyze the relationships between vertices and arcs of a graph. The adjacency matrix (or sociomatrix) of a directed graph $G=(V,E)$ with vertices $V=\{v_{1},\ldots,v_{n}\}$ is an $n\times n$ matrix $A_{G}$ where $a_{i,j}=1$ if $(v_{i},v_{j})\in E$, and 0 otherwise. For undirected graphs, this matrix is symmetric. In weighted graphs, entries $a_{i,j}$ can represent weights $w(v_{i},v_{j})$ instead of binary adjacency. In a simple directed graph, the incidence matrix $B_{G}$ with $n$ rows and $m$ columns captures relationships between nodes and arcs. Entries $b_{i,j}$ indicate whether node $v_{i}$ is the origin or destination of arc $e_{j}$. A simple graph is a graph without loops (self-edges) and multiple arcs between any pair of vertices. For a simple direct graph, we can assume by convention $b_{i,i}=0$ $\forall i\in I_{n}$, for $i\neq j$.
3 Overview of the Main Classical Centrality Measures
When examining relationships within a group, our initial inquiry often revolves around identifying influential individuals within the network. To address such queries effectively, it is essential to establish measures that quantify centrality. Many vertex centrality indices were introduced in the 1950s, such as the Bavelas index [2, 3], degree centrality [21], and a first feedback centrality introduced by Seeley [36]. Before providing a formal definition of centrality indices, it is necessary to introduce the concept of isomorphism.
Definition 1
Two directed graphs $G_{1}=(V_{1},E_{1})$ and $G_{2}=(V_{2},E_{2})$ are isomorphic (denoted as $G_{1}\simeq G_{2}$) if there exists a bijection $\varrho:V_{1}\rightarrow V_{2}$ such that for all $u,v\in V_{1}$:
$(u,v)\in E_{1}\Longleftrightarrow(\varrho(u),\varrho(v))\in E_{2}.$
Such a bijection is called an isomorphism.
An isomorphism that maps a graph onto itself is called an automorphism. Usually, we consider two graphs to be the same if they are isomorphic. Isomorphism and automorphism for a graph with symmetric relations are defined analogously.
Formally [24], a centrality index is a structural index which is defined as follows:
Definition 2 (Structural Index)
Let $G=(V,E)$ be a weighted, directed, or undirected multigraph, and let $V$ represent the set of vertices or edges of $G$, respectively. A real-valued function $s$ is called a structural index if and only if the following condition holds: for all $v\in V$, for any graph $H=(V^{\prime},E^{\prime})$ isomorphic to $G$, there exists a bijection $\varrho:V\rightarrow V^{\prime}$ such that $s_{G}(v)=s_{H}(\varrho(v))$, where $s_{G}(v)$ denotes the value of $s(v)$ in graph $G$.
A centrality index $C$ is required to be a structural index and thus induces at least a semi-order on the set of vertices or edges, respectively. According to this order, $v\in V$ is considered at least as central as $u\in V$ with respect to a given centrality $C$ if $C(v)\geq C(u)$. Centrality indices are structural measures used to quantify the importance of a node $v\in V_{G}$ in a graph $G$. For instance, $\sigma^{G}:V_{G}\rightarrow\Re_{0}^{+}$ assigns higher values to nodes that are more interconnected. In cases where two nodes $v$ and $u$ exhibit identical network structures, the centrality index should yield the same value $\sigma^{G}(v)=\sigma^{G}(u)$ for both nodes [28, 35]. Nodes $v$ and $u$ are considered identically structurally integrated into the network if there exists an automorphism $\eta:V_{G}\rightarrow V_{G}$ with $u=\eta(v)$. Centrality represents the extent to which an actor occupies a central position within a network. Empirical research has shown that occupying a central position in a network of relationships often correlates with access to resources and influence [14, 32, 33].
3.0.1 Degree
The number of edges connected to a vertex is referred to as its “degree” in graph theory. The degree of a node $v_{i}$ is commonly denoted as $d_{i}$. In the context of directed graphs, nodes have both in-degrees (edges pointing towards the node) and out-degrees (edges pointing away from the node), denoted as $d_{i}^{in}$ and $d_{i}^{out}$, respectively. For example, in social platforms like Facebook, the degree of a user corresponds to the number of friends [48].
Definition 3
In any undirected graph $G(V,E)$, the sum of all node degrees equals twice the number of edges:
$\sum_{i}d_{i}=2|E|$ (1)
This equation holds because each edge contributes to the degree sum of exactly two nodes.
In directed graphs, the sum of in-degrees equals the sum of out-degrees:
$\sum_{i}d_{i}^{in}=\sum_{j}d_{j}^{out}$ (2)
This equality arises because each directed edge contributes exactly one unit to the in-degree of its destination node and one unit to the out-degree of its origin node.
3.0.2 Degree centrality
In real-world interactions, individuals with many connections are often perceived as important. Degree centrality quantifies this idea into a measurable index. The degree centrality index ranks nodes higher based on the number of connections they have.
For an undirected graph, the degree centrality $C_{d}$ of a node $v_{i}$ is given by:
$C_{d}(v_{i})=d_{i}$ (3)
where $d_{i}$ represents the degree (number of adjacent edges) of node $v_{i}$.
In directed graphs, we distinguish between in-degree centrality, out-degree centrality, and total degree centrality, which combines in and out degrees:
[…]
In-degree centrality indicates how popular a node is within the network. Out-degree centrality measures how gregarious a node is.
Comparing degree centrality values across different networks directly is not straightforward due to variations in network size. To standardize comparison, degree centrality can be normalized as:
$C_{d}^{*}(v_{i})=\frac{d_{i}}{n-1}$ (7)
where $n$ is the number of vertices in the network.
3.0.3 Closeness centrality
Closeness centrality measures the average distance from a vertex to all other vertices in the network. Unlike other network metrics, closeness centrality offers a distinct perspective on each individual’s network position, capturing how quickly each vertex can reach all others [19]. The central idea is that nodes with higher closeness centrality are more central, allowing them to reach other nodes more efficiently [48]. Mathematically, closeness centrality $C_{c}(v_{i})$ for vertex $v_{i}$ is defined as:
$C_{c}(v_{i})=\frac{1}{\bar{l}_{D_{j}}}$ (8)
where
$\bar{l}_{D_{j}}=\frac{1}{n-1}\sum_{v_{j}\neq v_{i}}l_{i,j}$ (9)
represents the average shortest path length from node $v_{i}$ to all other nodes in the network. Nodes with smaller average shortest path lengths $l_{i,j}$ exhibit higher closeness centrality.
3.0.4 Betweenness centrality
Another perspective on centrality considers how crucial nodes are in connecting other nodes. For a node $v_{i}$, one approach is to quantify the number of shortest paths between other nodes that pass through $v_{i}$:
$C_{b}(v_{i})=\sum_{\begin{subarray}{c}s\neq v_{i}\neq t\end{subarray}}\frac{\sigma_{st}(v_{i})}{\sigma_{st}}$ (10)
Here, $\sigma_{st}$ denotes the total number of shortest paths from node $s$ to $t$ (also referred to as information pathways), and $\sigma_{st}(v_{i})$ represents the number of those paths that pass through $v_{i}$. Essentially, this metric evaluates how central $v_{i}$ is in facilitating connections between any pair of nodes $s$ and $t$, known as betweenness centrality.
To standardize betweenness centrality across networks, normalization is necessary. This involves calculating its maximum possible value:
$C_{b}^{*}(v_{i})=\frac{C_{b}(v_{i})}{\max_{s\neq t}C_{b}(v_{i})}$
The maximum value occurs when $v_{i}$ lies on all shortest paths from $s$ to $t$ for every pair $(s,t)$, i.e., $\forall(s,t),s\neq t\neq v_{i},\frac{\sigma_{st}(v_{i})}{\sigma_{st}}=1$.
4 A brief introduction to Fuzzy Numbers
The limit of bivalent logic is that it addresses the real problems as if they were white or black [25]. Understanding the critical point at which a thing becomes a non-thing undermines traditional logic, which is based on black and white, all or nothing, true or false; according to Aristotelian logic, an element belongs to a set or does not belong at all. Aristotelian thought is essentially based on the principles of identity, non-contradiction, and the excluded middle. Since ancient times, Mathematicians and Philosophers have tried to find a rational explanation for paradoxes to save the dichotomy assumption and avoid dealing with the grey that exists in all things. The first to use the term fuzzy to describe logic based on uncertainty and vagueness was Lotfi Zadeh. When Zadeh published his first article on “Fuzzy Sets”, he introduced the concept of “membership value” which, in contrast to the bivalent theory, considers the possibility of intermediate values of truth; in particular, the degree of membership can take any value in the range $[0,1]$[45]. Fuzzy logic contemplates the grey and starts where the contradictions inherent in things begin, where A and non-A melt to a certain extent. Thus, fuzzy logic essentially deals with uncertainty due to vagueness.
In classical set theory, a set comprises all the elements of the discourse universe that satisfy a given membership function. For a traditional set, which we define as crisp, to distinguish it from a fuzzy set, the membership function is Boolean; it associates to each element $x$ of the universe a value of “true” or “false” depending on whether or not $x$ belongs to the set. Formally, a crisp set has the following characteristic function:
$\mu_{A}(x)=\begin{cases}1&\text{if x $\in$ A}\\ 0&\text{if x $\notin$ A}\end{cases}$ (11)
The theory of fuzzy sets posits that an element can partially belong to a set, determined by a membership function that assigns real values in the interval $[0,1]$.

### v2

# Novel Fuzzy Centrality Measures in Vague Social Networks

@@ Abstract
Social network analysis (SNA) helps us understand the relationships and interactions between individuals, groups, organizations, or other social entities. In the literature, ties are generally considered binary or weighted based on their strength. Nonetheless, when the actors are individuals, these relationships are often imprecise, and identifying them with simple scalars leads to information loss. Indeed, social relationships are often vague in real life, and although previous research has proposed the use of fuzzy networks, these are typically characterized by crisp ties. The use of weighted links does not align with the original philosophy of fuzzy logic, which instead aims to preserve the vagueness inherent in human language and real life. For this reason, this paper proposes a generalization of the so-called Fuzzy Social Network Analysis (FSNA) to the context of imprecise relationships among actors. Dealing with imprecise ties and introducing fuzziness in the definition of relationships requires an extension of social network analysis, defining ties as fuzzy numbers instead of crisp values and extending classical centrality indices to fuzzy centrality indexes. The article presents the theory and application of real data collected through a fascinating mouse-tracking technique to study the fuzzy relationships in a collaboration network among the members of a university department.
Keywords: Fuzzy Social Network, Fuzzy Centrality Indices, Vague Relationships, Uncertainty, Fuzzy Numbers, University Department Collaboration Network.
@@ 1 Introduction
Social network analysis (SNA) helps us understand the relationships and interactions between individuals, groups, organisations, or other social entities [Hanneman and Riddle, 2011, Otte and Rousseau, 2002, Opsahl et al., 2010, Wasserman, 1994, Landherr et al., 2010]. Each type of social structure can be represented with a graph, in which the social entities are represented as nodes or vertices in a network, and their relationships are represented as edges or ties. SNA examines relationships between entities and how information, resources, or influence flows through them. It identifies patterns, structures, and dynamics of social systems. Ties in a network are generally binary based on the presence or absence of the link or weighted based on their strength. Each connection can be identified by a number in $[0,1]$ or $\{0,1\}$, respectively.
Nonetheless, when the actors of a social network are individuals, the relationships are very complex and measuring them using a single number leads to a loss of information. In other words, relationships are inherently imprecise in real life, and thus, social network analysis should consider vagueness. Consider a social network like Facebook, where nodes represent users and links denote friendships. However, not all friendships carry the same meaning: some are more interactive, while others are more distant. This mirrors the complexity of real-life relationships, which are rarely clear-cut but often vague. The concept of "friendship" or "collaboration" is inherently ambiguous, varying with context, and highlighting the uncertainty present in both online interactions and real-life connections. The concepts of friendship, close friendship, collaborations, or acquaintances are inherently vague, and representing them with a single number sacrifices information for the sake of simplicity.
The most natural way to consider the problem mentioned above is to introduce a fuzzy-based social network analysis. In the literature, fuzzy network analysis has mainly been studied by introducing weighted ties within a network [see, e.g., Hu et al., 2015, Brunelli and Fedrizzi, 2009]. However, weighting the links between nodes has nothing to do with fuzzy logic, and using the term “fuzzy” raises some doubts because it suggests an approach that considers the vagueness of relationships, which only in this latter case would be in line with the basic idea of fuzzy logic that begins and thrives where opposites coexist. Indeed, fuzzy logic deals with situations where A and non-A coexist simultaneously with a certain degree of truth [Zadeh, 1965, 1975, 1968, Yager, 1980]. Hence, translating vague concepts into crisp values necessarily results in a loss of information regarding the inherently imprecise nature of these expressions. In other words, assigning a link a value in $[0,1]$ or $\{0,1\}$ can sometimes be a simplification that sacrifices information for the sake of convenience.
Dealing with imprecise ties and introducing fuzziness in the definition of relationships needs an extension of classical social network analysis to fuzzy numbers instead of crisp values. The mathematical formalisation of this generalisation requires rewriting the classical centrality indices into fuzzy centrality indices involving fuzzy relationships. The main motivation of this study is that, in many contexts, relationships between nodes in a social network should be characterised by functions rather than scalars to preserve the information about the vagueness of ties. The article proposes new fascinating Fuzzy Centrality Measures (FCMs) for Fuzzy Social Networks (FSNs) and show an application to real data. The dataset adopted in this study is collected through a mouse tracking technique to capture the vagueness of the relationships among the members of a University department. The best results of this research and the main advantages of our proposal are the creation of social networks based on vague ties and centrality indices preserving information about their imprecision, inherent in human nature, human language, and interpersonal relationships.
@@
The first part of the work introduces the preliminaries necessary for formally understanding the proposed indices. Subsequently, based on the idea mentioned above, new centrality indices are proposed: fuzzy degree centrality, fuzzy out-degree centrality index, total fuzzy degree centrality index, fuzzy betweenness centrality, fuzzy closeness centrality, and fuzzy out-closeness centrality. The application Section proposes a fascinating investigation based on a University department collaboration network. The paper ends with a discussion and conclusions.
2 Background on Centrality Indices in Social Networks
Centrality indices are essential tools in social network analysis, capturing the importance and influence of nodes within a network, and have been a topic of research for many decades. Bavelas [1948] pioneered centrality measures in the paper “A Mathematical Model for Group Structures” presenting one of the earliest formal approaches to understanding social networks and group structures through mathematical modeling. He explores how the arrangement of communication pathways within a group affects the group’s efficiency, centralization, and overall functioning. Katz [1953] introduced the concept of Katz centrality, which extends the idea of degree centrality by considering the influence of both direct and indirect connections. Katz centrality assigns more weight to closer (direct) connections while also accounting for longer paths in the network. The concept of centrality has been explored in various forms, with foundational contributions such as the general framework for centrality indices proposed by Sabidussi [1966], which provided a unified approach to understanding node importance within a network. Closeness centrality was introduced by Linton C. Freeman in his seminal 1979 paper titled “Centrality in Social Networks: Conceptual Clarification[Freeman, 2002]. In this paper, Freeman formalized various centrality measures, including closeness centrality, which measures how close a node is to all other nodes in a network. The concept was developed to capture the idea that a node is more central if it can quickly interact with all other nodes, reflecting its potential to communicate or spread information efficiently within the network. In their book “Social Network Analysis: Methods and Applications”, Wasserman [1994] provided a comprehensive treatment of various centrality measures and their applications in social network analysis. Their work has been instrumental in standardizing centrality concepts and methods in the field. Opsahl et al. [2010] proposed generalizations of degree and shortest-path centralities to weighted networks, providing a more comprehensive view of centrality in networks where connections have varying strengths. These measures take into account both the number and weight of connections, offering a more accurate reflection of a node’s importance. These indices remain fundamental in understanding the roles and significance of nodes within complex networks. In their critical review, Landherr et al. assessed various centrality measures, highlighting the strengths and limitations of each within different network contexts. This work is valuable for understanding how different centrality measures can be applied depending on the specific characteristics of a network [Landherr et al., 2010]. Blue et al. [2002] contributed to the application of fuzzy logic in graph theory and, by extension, social networks. Their work involves developing fuzzy graph models that can be applied to social network analysis, allowing for the representation of uncertainty and partial membership in social ties. Brunelli and Fedrizzi [2009] discussed a fuzzy approach to social network analysis, including the development of fuzzy centrality indices. Their approach considered the fuzziness of social ties and how this impacts the calculation of centrality within networks. However, in their framework, the ties between nodes are represented by crisp membership values, which indicate the degree of membership or connection strength within the fuzzy context.
3 Preliminaries
3.1 Fuzzy Numbers